Circular No. 8050

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1720–318 R. A. Remillard, A. M. Levine, and E. H. Morgan, Massachusetts

Institute of Technology (MIT); and E. Smith and J. Swank, Goddard Space Flight Center (GSFC), for the RXTE All Sky Monitor Team at MIT and NASA’s GSFC, report the discovery of a transient x-ray source at α = 17^{h}19^{m}58^{s}, δ = −31^{o}46′.8 (equinox 2000.0; estimated 3′ uncertainty at 90percent confidence): “The source was first detected on 2003 Jan. 9, when the average flux (2–12 keV) was 130 ± 20 mCrab. The intensity increased to 430 ± 40 mCrab on Jan. 10, and to 400 ± 20 mCrab on Jan. 13. A brief observation of the ASM position on Jan. 14 with the RXTE PCA confirms the presence of a bright transient. More extensive observations with the PCA and HEXTE, including raster scans to improve the source position, are planned starting Jan. 15. The error circle does not contain any noteworthy sources in the Simbad catalogs. The ASM hardness ratios indicate that the spectrum was moderately hard during Jan. 9–10 and relatively soft on Jan.

13. These results resemble the early spectral evolution of x-ray transients known to contain an accreting black hole. We encourage optical and radio observations of this new transient.”

COMET C/2002 V1 (NEAT) M. L. Sitko, University of Cincinnati; and D. K. Lynch, R. W. Russell,

and D. Kim, The Aerospace Corporation, report that 3–14-μm spectroscopy of C/2002 V1 (NEAT), obtained on Jan. 9.2 UT with the Aerospace Broadband Array Spectrograph System at the NASA Infrared Telescope Facility 3-m reflector, revealed a thermal-emission spectrum showing a trapezoidalshaped silicate emission feature with shoulders at 9.0 and 11.2 μm. The underlying continuum was well fitted with a 290-K black body, ≈ 14 percent above the black-body radiative equilibrium temperature of 254 K. Narrowband magnitudes (± 0.1) in a 3′′.5-diameter aperture were M [4.5 μm] = 8.5 and N [10.2 μm] = 3.1. On the following night, the comet was ∼ 10 percent brighter, but otherwise unchanged.

SUPERNOVA 2003I IN IC 2481 T. Kobayashi, Oizumi, Gunma-ken, Japan, reports (via S. Nakano,

Sumoto, Japan) that SN 2003I appeared at mag 17.8 on an unfiltered CCD prediscovery image taken on Jan. 8.604 UT with a 0.41-m f/4.3 reflector.

2003 January 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green